OpenStudy (anonymous):

can someone explain how to find the limit for this function: lim as deltax approaches 0 for 2(x+deltax)- 2x all over deltax

OpenStudy (aum):

\[ \lim_{\Delta x \rightarrow 0}\dfrac{2(x+\Delta x) - 2x}{\Delta x} \]

OpenStudy (aum):

Simplify the numerator.

OpenStudy (aum):

\[ \lim_{\Delta x \rightarrow 0}\dfrac{2(x+\Delta x) - 2x}{\Delta x} = \lim_{\Delta x \rightarrow 0}\dfrac{2x+2\Delta x - 2x}{\Delta x} = 2 \]

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thank you!!

OpenStudy (aum):

you are welcome.

OpenStudy (anonymous):

but wait i'm still kind of lost wouldnt it be undefined since both delta x equals 0 ?

OpenStudy (aum):

They cancel each other out.

OpenStudy (cwrw238):

2 * delta x / delta x = 2

OpenStudy (aum):

Note: When taking limit as delta x approaches zero, remember the delta x does NOT actually become zero. You find the limit as delta x APPROACHES zero. So you are not really dividing zero by zero.

OpenStudy (anonymous):

oh okay, i see what you're saying. so the limit would equal to 2. i get it now!

OpenStudy (aum):

Yes.